• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 16:06:25 on December 7,2006

    Entry number 190267

    keyword=Shift Summary

    8:00 AM Day shift begins.
    R. Lindgren SL
    O. Hansen TO
    Planned stop of experiment at 8:00 AM
    mCC optimizing polarized source to get more beam current. Want 70 ua.
    Hansen was relieved of duties and Lindgren was both SL and TO.
    0 hours of beam during the day shift.
    Cryo-target parked in dummy cell position all day.
    No target problems.
    During the day Hall A was in controll access.
    Work was being dome on installing a new trigger for helium3 recoil arm
    and on repairing a bad cryo-level meter in the right HRS.
    MCC still working on source.
    16:00 Shift Ends